OpenStudy (anonymous):

given F(x)=\sqrt{(3x-9)}^{3} find f ^{-1}(x)

OpenStudy (freckles):

\[f(x)=(\sqrt{3x-9})^3 \\ y=(\sqrt{3x-9})^3\] solve for x

OpenStudy (anonymous):

how do you solve for x

OpenStudy (freckles):

well first is what I wrote what you mean?

OpenStudy (anonymous):

well the first one is but then i have to find f^-1(x)

OpenStudy (freckles):

ok I just wanted to make sure I wrote your function correctly

OpenStudy (freckles):

try raising both sides to (1/3)

OpenStudy (freckles):

we are doing this since 3*1/3 will give us 1

OpenStudy (anonymous):

ok so what do we do with the 1

OpenStudy (freckles):

anything to the first power can be written without the exponent that is u^1=u

OpenStudy (anonymous):

okay

OpenStudy (freckles):

when you are ready you can show me what you have after doing that?

OpenStudy (anonymous):

\[f^-1x\]

OpenStudy (anonymous):

thats supposed to be srrt -1x

OpenStudy (freckles):

\[y=(\sqrt{3x-9})^3\] so asked you to raise both sides to (1/3) so you should have had this: \[y^\frac{1}{3}=(\sqrt{3x-9})^{3 \frac{1}{3}} \\ y^\frac{1}{3}=(\sqrt{3x-9})^1 \\ y^\frac{1}{3}=\sqrt{3x-9}\] but I don't even you tried doing that based on what you have there :(

OpenStudy (freckles):

anyways square both sides not so you can undo that radical part that contains the x thingy majigger

OpenStudy (anonymous):

i thought you meant replace the square root im sorry

OpenStudy (freckles):

square both sides so you*

OpenStudy (freckles):

no I didn't say that I said to raise both sides to (1/3) power

OpenStudy (anonymous):

yeah i misunderstood

OpenStudy (freckles):

ok do you want to try squaring both sides ?

OpenStudy (freckles):

another way to say that is raise both sides to second power (we are trying to undo the square root thing)

OpenStudy (anonymous):

okay i have to log off but i get what your saying and i understand it so thank you

OpenStudy (freckles):

ok cool your main objective is to solve for x and then you can interchange the variables at the end

OpenStudy (freckles):

but later and good luck @kimdarlene

OpenStudy (anonymous):

thank you and you to